OpenStudy (anonymous):

what is an equation of the line that passes through the points (-4,3) and (-1,6) A. y=-x-7 B. y=-x-1 C. y=7x+1 D. y=x+7

OpenStudy (anonymous):

do you know how to find the slope ?

OpenStudy (anonymous):

is it the y2-y1 thing ?

OpenStudy (anonymous):

slope(m) = (y2 - y1) / (x2 - x1) (-4,3) x1 = -4 and y1 = 3 (-1,6) x2 = -1 and y2 = 6 now we sub slope(m) = (6 - 3) / (-1 - (-4) m = 3/(-1 + 4) m = 3/3 m = 1 <-- your slope understand me so far ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

which answer choice is it ?

OpenStudy (anonymous):

now we can use y = mx + b, because we know the slope is 1 and we need to find the y intercept, which is b. y = mx + b m = 1 (-1,6) x = -1 and y = 6 sub y = mx + b 6 = 1(-1) + b 6 = -1 + b 6 + 1 = b 7 = b making your equation : y = x + 7

OpenStudy (anonymous):

.thank you so much

OpenStudy (anonymous):

no problem :)